# Unlearning vs. Obfuscation: Are We Truly Removing Knowledge?

### v1

@@ Abstract
Unlearning has emerged as a critical capability for large language models (LLMs) to support data privacy, regulatory compliance, and ethical AI deployment. Recent techniques often rely on obfuscation by injecting incorrect or irrelevant information to suppress knowledge. Such methods effectively constitute knowledge addition rather than true removal, often leaving models vulnerable to probing. In this paper, we formally distinguish unlearning from obfuscation and introduce a probing-based evaluation framework to assess whether existing approaches genuinely remove targeted information. Moreover, we propose DF-MCQ, a novel unlearning method that flattens the model predictive distribution over automatically generated multiple-choice questions using KL-divergence, effectively removing knowledge about target individuals and triggering appropriate refusal behaviour. Experimental results demonstrate that DF-MCQ achieves unlearning with over 90% refusal rate and a random choice-level uncertainty that is much higher than obfuscation on probing questions. [^1] https://github.com/potsawee/unlearning-hallu/tree/mcq
@@ 2 Related Work
Obfuscation-based methods introduce misleading or confusing information into the training data to obscure learned knowledge (i.e., a form of knowledge addition), thus indirectly causing forgetting. Notable techniques such as WHP Eldan and Russinovich (2023) and WHP+ Liu et al. (2024) achieve knowledge removal by overwhelming models with conflicting knowledge, thus reducing model confidence in previously learned facts. UnStar Sinha et al. (2025) further develops this approach by using better counter samples with misleading rationales, disrupting original knowledge. Although obfuscation-based methods are effective at preventing access to the original information, they mask rather than fully erase knowledge, rendering them susceptible to leakage under carefully designed probing conditions Xu et al. (2025b).
@@ 3 Unlearning and Obfuscation
We consider unlearning from an uncertainty perspective by treating the entire model knowledge as a knowledge graph. Given a forget set =$\mathcal{F}=\{(X_{i},R_{i}^{j},Y_{i}^{j})\}_{i=1}^{N}$ containing a number of facts $F_{i}^{j}$ (i.e., triplets) that should be removed. Each fact contains a subject $X_{i}$ (e.g., Wilhelm Wattenbach), its relevant object $Y_{i}^{j}$ (e.g., Rantzau), which is connected by the relations $R_{i}^{j}$ (e.g., born in). Let $\theta$ be the model parameters, we define the unlearning effect as follows:
$H_{\theta}(Y_{i}|X_{i},R_{i}^{j};\mathcal{D})\approx H_{\theta}(Y_{i}|X_{i},R_{i}^{j};\mathcal{D}\textbackslash F_{i}^{j})$ (1)
where $Y_{i}\in\mathcal{Y}_{i}$ represents all possible objects following $X_{i}$ and $R_{i}^{j}$, $\mathcal{D}$ represents the training data of the LLM and $\mathcal{D}\textbackslash F_{i}^{j}$ is the training data excluding the fact $F_{i}^{j}$. The entropy $H_{\theta}(Y_{i})=-\sum_{Y_{i}\in\mathcal{Y}_{i}}P(Y_{i})\log P(Y_{i})$. That is, the model has the same level of uncertainty as one that is trained on the dataset excluding fact $F_{i}^{j}$. For a non-hallucinatory instruction-tuned LLM nowadays, when prompted with a query it does not have an answer to, the model will refuse to answer or explicitly indicate that it does not have the knowledge. Therefore, an indication of unlearning effect is the model refusal behaviour, as follows.
@@ 7.1 Open-Ended Questions
Table 2: Performance comparison among NPO, WHP+ and DF-MCQ on WPU using Llama-3.1-8B-Instruct. Forget set, retain set and hard retain set performance are measured by ROUGE-L recall. The refusal rate is the percentage of responses that refuses to answer questions in the forget set.
@@
Table 3: WHP+ and DF-MCQ on WPU using Qwen2.5-7B-Instruct. Forget set and hard retain set performance are measured by ROUGE-L recall. The refusal rate is the percentage of responses that refuses to answer questions in the forget set.
@@ 7.2 Yes-No Probing Questions
Figure 5: The rate of answering Yes (Yes rate) against learning rates and number of obfuscation samples (a) and correlation between unlearning efficacy and Yes rate (b). Each point in (b) corresponds to a point in (a) with unlearning efficacy measured by $1-$ROUGE-L. The Pearson Correlation Coefficient of (b) is 0.84.
@@
First, increasing the number of obfuscation samples increases the tendency of shortcut. Second, with an increasing learning rate and hence larger model updates, the model is more likely to answer Yes. We also plot the correlation between unlearning efficacy measured by $1-$ROUGE-L on the open-ended questions and the Yes rate as shown in Fig. 5 (b). The Pearson Correlation Coefficient is 0.84. That is, to obfuscate the model to a degree that effectively protects privacy, the model is very likely to answer Yes to all probing questions about the target person.
@@ Appendix A Break Down Results for MCQ Probing
We provide breakdown results for MCQ probing questions to show the possible failure mode of obfuscation on specific individuals. The performance of the original model, WHP+ and DF-MCQ are shown in Tables 6, 7 and 8 respectively.

### v2

@@ Abstract
Unlearning has emerged as a critical capability for large language models (LLMs) to support data privacy, regulatory compliance, and ethical AI deployment. Recent techniques often rely on obfuscation by injecting incorrect or irrelevant information to suppress knowledge. Such methods effectively constitute knowledge addition rather than true removal, often leaving models vulnerable to probing. In this paper, we formally distinguish unlearning from obfuscation and introduce a probing-based evaluation framework to assess whether existing approaches genuinely remove targeted information. Moreover, we propose DF-MCQ, a novel unlearning method that flattens the model predictive distribution over automatically generated multiple-choice questions using KL-divergence, effectively removing knowledge about target individuals and triggering appropriate refusal behaviour. Experimental results demonstrate that DF-MCQ achieves unlearning with over 90% refusal rate and a random choice-level uncertainty that is much higher than obfuscation on probing questions. [^1] https://github.com/potsawee/unlearning-dfmcq
@@ 2 Related Work
Obfuscation-based methods introduce misleading or confusing information into the training data to obscure learned knowledge (i.e., a form of knowledge addition), thus indirectly causing forgetting. Notable techniques such as WHP Eldan and Russinovich (2023) and WHP+ Liu et al. (2024) achieve knowledge removal by overwhelming models with conflicting knowledge, thus reducing model confidence in previously learned facts. UnStar Sinha et al. (2025) further develops this approach by using better counter samples with misleading rationales, disrupting original knowledge. FLAT Wang et al. (2025) maximizes difference between their designed template answer and forget answer to avoid using a retain set, and RKLD Wang et al. (2024) decreases the probability of the most likely option while increase the probability of runner-ups. Although obfuscation-based methods are effective at preventing access to the original information, they mask rather than fully erase knowledge, rendering them susceptible to leakage under carefully designed probing conditions Xu et al. (2025b). Moreover, Hu et al. (2025) argues that existing unlearning methods merely obscure the target information, as shown by their success with a fine-tuning attack. In contrast, this work draws a distinction between obfuscation and unlearning, and we directly probe unlearned models, without additional fine-tuning.
@@ 3 Unlearning and Obfuscation
We consider unlearning from an uncertainty perspective by treating the entire model knowledge as a knowledge graph. Given a forget set $\mathcal{F}=\{(X_{i},R_{i}^{j},Y_{i}^{j})\}_{i=1}^{N}$ containing a number of facts $F_{i}^{j}$ (i.e., triplets) that should be removed. Each fact contains a subject $X_{i}$ (e.g., Wilhelm Wattenbach), its relevant object $Y_{i}^{j}$ (e.g., Rantzau), which is connected by the relations $R_{i}^{j}$ (e.g., born in). Let $\theta$ be the model parameters, we define the unlearning effect as follows:
[…]
@@ 7.1 Open-Ended Questions
Table 2: Performance comparison of NPO, WHP+ and DF-MCQ on open-ended questions from WPU using Llama-3.1-8B-Instruct. Forget set, retain set and hard retain set performance are measured by ROUGE-L recall. The refusal rate is the percentage of responses that refuses to answer questions in the forget set.
@@
Table 3: WHP+ and DF-MCQ on open-ended questions from WPU using Qwen2.5-7B-Instruct. Forget set and hard retain set performance are measured by ROUGE-L recall. The refusal rate is the percentage of responses that refuses to answer questions in the forget set.
@@
Robustness to SFT: We show the robustness of models unlearnt with DF-MCQ to the supervised fine-tuning (SFT) attack with questions about other individuals in Table 4. As a result, SFT attack does not influence the forget set performance, and the ones that output "I don’t know" still outputs "I don’t know" after finetuning.
Model Forget Hard Retain
DF-MCQ 10.70 60.53
+ SFT Attack Hu et al. (2025) 10.70 61.33
Table 4: Robustness to SFT attack on WPU test set.
Continual Unlearning: The DF-MCQ can be applied as a continual unlearning method that we can continue adding new unlearning targets without affecting previously unlearnt targets. In theory, this method can work for any number of targets. The results of continually unlearn the 10 individuals, compared to the performance of unlearn just a pair, are shown in Table 5.
Model Forget Hard Retain Refusal
DF-MCQ Forget 2 10.70 60.53 92.72
DF-MCQ Forget 10 11.70 58.77 91.81
Table 5: Continually unlearn 10 (Forget 10) individuals comapred to the average performance of the standard unlearn 2 setting in Table 2 (Forget 2).
@@ 7.2 Yes-No Probing Questions
Figure 5: The rate of answering Yes (Yes rate) against learning rates and number of obfuscation samples (a) and correlation between unlearning efficacy and Yes rate (b). Each point in (b) corresponds to a point in (a) with unlearning efficacy measured by 11-ROUGE-L. The Pearson Correlation Coefficient of (b) is 0.84.
@@
First, increasing the number of obfuscation samples increases the tendency of shortcut. Second, with an increasing learning rate and hence larger model updates, the model is more likely to answer Yes. We also plot the correlation between unlearning efficacy measured by 11-ROUGE-L on the open-ended questions and the Yes rate as shown in Fig. 5 (b). The Pearson Correlation Coefficient is 0.84. That is, to obfuscate the model to a degree that effectively protects privacy, the model is very likely to answer Yes to all probing questions about the target person.
@@ Appendix A Break Down Results for MCQ Probing
We provide breakdown results for MCQ probing questions to show the possible failure mode of obfuscation on specific individuals. The performance of the original model, WHP+ and DF-MCQ are shown in Tables 8, 9 and 10 respectively.